OpenStudy (anonymous):

Please help! I've been working on this for two hours! Simplify this expression and state any restrictions on the variables: (x+1)/(x^2+2x-8) - (x/(4x-8)

OpenStudy (anonymous):

@amistre64 @Agent_Sniffles @TuringTest

OpenStudy (anonymous):

Can you factor a number out of that second term?

OpenStudy (anonymous):

Can't you factor out a two?

OpenStudy (anonymous):

You can factor out a four too.

OpenStudy (anonymous):

Okay, so how do I go about doing that?

OpenStudy (campbell_st):

well factorise both denominators \[\frac{x +1}{(x +4)(x -2)} - \frac{x}{4(x -2)}\] you'll need a common denominator of 4(x + 4)(x -2) so the 1st fraction is multiplied by 4/4 and the 2nd fraction is multiplied by (x +4)/(x +4) which will give \[\frac{4(x + 1)}{4(x +4)(x -2)} - \frac{x(x +4)}{4(x +4)(x -2)}\]

OpenStudy (anonymous):

hehe like so! ^

OpenStudy (anonymous):

Sorry - I didn't mean you could factor out a four AND a two. You can factor out a two or a four. A four is a better idea though

OpenStudy (campbell_st):

the restrictions are found by looking at the denominator and solving x + 4 = 0 x - 2 = 0 just simplify the numerator

OpenStudy (anonymous):

so is the answer ((4x+4)/(x+4)) - ((x^2+4x)/(x-2))?

OpenStudy (anonymous):

How did you cancel the 4(x-2) on the left term and the 4(x+4) term on the right? I don't think its possible.

OpenStudy (anonymous):

I'm not sure.... What should the answer be?? I'm confused.. Sorry, I just need a good grade on this

OpenStudy (campbell_st):

well start by leaving the denominator alone \[\frac{ 4x + 4 - x^2 - 4x}{4(x+4)(x -2)}\] see if you can simplify the numerator

OpenStudy (anonymous):

(-8x^2-x^2+4)/((4(x+4)(x-2)) ??

OpenStudy (campbell_st):

so the numerator can be written as \[\frac{-(x^2 -4)}{4(x + 4)(x -2)}\] the numerator can be factorised... its the difference of 2 squares... a common factor can be removed. but a point of discontinuity will exist...

OpenStudy (anonymous):

ok so the answer is (-(x^2-4)/4(x+4)(x-2)) and no restrictions?

OpenStudy (anonymous):

Try factoring the numerator. There are restrictions.

OpenStudy (campbell_st):

there are restrictions... 1 a vertical asymptote 2. a point of discontinuity